The equity shares of a firm in the current stock market has been traded at Rs60 per share. The price earnings ratio is 10 times. The Dividend payout ratio is 75%. The total number of shares issued and outstanding as on date are 100000 equity shares of Rs10 each. Book value of each share is Rs40 Describe and Compute - Earnings per share, return on equity.


1
Expert's answer
2021-10-18T11:29:29-0400

Earning per share is the company's profit divided by the company's outstanding shares.


"EPS = \frac{Net profit}{Outstanding shares}"


where "EPS -" is the earnings per share


The price earnings ratio "\frac{P}{E}= \frac{Share price}{Earnings per share}"


To determine EPS


"\frac{P}{E} is provided = 10"


Share price "= 60"


Therefore EPS "= \frac{share price}{Price earning ratio}"


"EPS =\frac {60}{10} = 6"



RETURN ON EQUITY


Return on equity shows investors how efficiently a company is handling the investors' money.


Return on Equity ROE "= \frac{Net Income}{Shareholders Equity}"


To determine ROE


Dividend Payout Ratio "= \frac{Total Dividend}{Net income}"


but Dividend payout ratio "= 75\%"


Total dividend = "= Total equity * earning per share"


Total divided "= 100,000 * 6 = 600,000"


Therefore "\therefore"

Net income "= \frac{Total divided}{Divided payout ratio}"


Net income "= \frac{600000}{75\%} = 450,000"


Shareholders equity "=Total number of shares * book value of each share"


Shareholders equity "= 100,000 * 40 = 4,000,000"

"\therefore"


"ROE = \frac{450,000}{4000,000} *100 = 11.25\%"